OpenStudy (cherie_magee):

In a certain clothing store, 6 shirts and 3 ties cost $79.50, and 3 shirts and 2 ties cost $41. Determine the cost of each tie. $2.50 $3.50 $4.50

OpenStudy (anonymous):

your job is the unenviable task of solving \[6x+3y=79.50\\3x+2y=41\]

OpenStudy (anonymous):

damn that was wrong lets try again multiply the second one by \(-2\) and get \[6x+3y=79.50\\-6x-4y=-82\] add and get \[-y=-2.50\] making \(y=2.50\)